# Triple and binary evolution pathways play equally significant roles in LISA double white dwarf formation

### v1

@@ 2.2 Initial conditions
Figure 1: Initial parameter distributions. The left panel shows the mass distributions, where $m_{1}$ and $m_{2}$ denote the masses of the inner binary components, and $m_{3}$ represents the mass of the tertiary. The middle panel displays the semi-major axis distributions for the inner ($a_{1}$) and outer ($a_{2}$) orbits, respectively. The right panel illustrates the eccentricity distributions of the inner ($e_{1}$) and outer ($e_{2}$) orbits, respectively. See Section 2.2 for more details.
@@
We draw the primary mass $m_{1}$ from a Kroupa initial mass function (Kroupa, 2001) between $1$ and $8\,\mathrm{M_{\odot}}$. Furthermore, we follow functions from Moe & Di Stefano 2017 to sample the orbital period ($0.2\leq\log({T}_{1}/\text{days})\leq 8$) and secondary mass ( $0.08\,\mathrm{M_{\odot}}\leq m_{2}\leq m_{1}$) of the inner binary, and calculate the semi-major axis $a_{1}$ from Kepler’s law. Similarly, the orbital period of the outer binary also follows Moe & Di Stefano 2017, where we assume that the inner binary is represented as a single star with a mass of $m_{1}+m_{2}$. We allow the tertiary mass $m_{3}$ to be more massive than the total mass of the inner binary in certain cases and an extrapolated mass ratio distribution from Moe & Di Stefano 2017 is used to sample $m_{3}$. In addition, we sample the eccentricities of inner ($e_{1}$) and outer ($e_{2}$) orbit from Moe & Di Stefano 2017 and randomly sample the spatial orientations of the inner and outer orbital frames from isotropic distributions.
We reject any star whose radius exceeds the Roche lobe radius on the zero-age main sequence (Eggleton, 1983) and reject any system which would be dynamically unstable (Vynatheya et al., 2022) at the start of the simulation. Any Roche lobe overflow or dynamical instabilities are modeled using prescriptions in MSE during the evolution (see Section 2.1 for more details).
@@ 2.3 Construction of Galactic double white dwarf population
We use a Milky Way-like galaxy from the large-scale cosmological magneto-hydrodynamical simulations TNG50 (Nelson et al., 2019; Pillepich et al., 2019). With a comoving volume of $(50\,\mathrm{Mpc})^{3}$ the TNG50 simulation box contains about 100 Milky Way-like galaxies with a total mass of $10^{14}\,\mathrm{M_{\odot}}$. To select a suitable Milky Way-like galaxy, we randomly choose one of the six halos with a total mass $1$ – $2\times 10^{12}\,\mathrm{M_{\odot}}$ and whose central galaxy has a stellar mass $5$ – $7\times 10^{10}\,\mathrm{M_{\odot}}$. Additionally, we examine the galaxy’s stellar projection to confirm disk dominance. The mass of the selected galaxy (Galaxy ID = 476266) is $\sim 5\times 10^{10}\,\mathrm{M_{\odot}}$ which is consistent with our Milky way galaxy (e.g. Bland-Hawthorn & Gerhard, 2016). We extract present-day properties such as age, stellar mass, and 3D-position of the star particles. We place our observer at a randomly assigned Sun-like position in the disk, $8.2\,\rm kpc$ from the Galactic center. We then measure the distances to all LISA-detectable double white dwarfs from each star particle to this location. We combine these Galactic properties with the simulated triples to construct a representative Galactic double white dwarf population as follows.
From the $10^{5}$ simulated triples with MSE, we select those which form double white dwarfs that emit gravitational waves in the LISA frequency bandwidth ($10^{-4}\,\mathrm{Hz}$ – 0.1 Hz) during their evolution. We use this sample to assemble the entire Galactic LISA double white dwarf population in each star particle of the selected TNG50 Milky Way-like galaxy. In our simulated sample all the stars are assumed to be formed simultaneously. To allow for continuous star formation we seed the number of LISA double white dwarfs per star particle ($N_{\textrm{DWD},\star}$) based on the age and stellar mass ($M_{\star}$) of the star particle as:
$N_{\textrm{DWD},\star}=\frac{N_{\textrm{DWD, MSE}}}{M_{\textrm{tot, MSE}}}\times M_{\star},$ (1)
where $N_{\text{DWD, MSE}}$ is the total number of all double white dwarfs in MSE which emit gravitational waves in the frequency bandwidth of LISA at a simulation time equal to the age of the star particle and $M_{\text{tot, star\, particle}}$ is the mass of the star particle at $z=0$, which is obtained from the TNG50 simulation. The total stellar mass of the simulated population is given by
@@
for $\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{for }f_{t}\neq 0,$
@@ 3 Key processes that shape the triple evolutionary pathways
In this section, we present a brief overview of the five key processes characteristic of triple evolution that shape the evolutionary pathways leading to the formation of LISA double white dwarfs:
@@
    Triple common envelope: The third star initiates mass transfer onto the inner binary, leading to the formation of a LISA double white dwarf.
@@
We show a schematic diagram for primary processes that drive evolutionary pathways leading to LISA double white dwarfs from triples in Figure 2. In the following subsections, we discuss these processes in greater depth with detailed examples. All error bars presented below are estimated by scaling up the fractional Poisson error from the intrinsic population evolved with MSE. Figure 3 shows that the processes are not mutually exclusive. For instance, about 4.8% of Galactic LISA double white dwarfs from triples undergo a TCE phase and induced mass transfer, leading to a merger in the inner binary. Hence, the relative percentages quoted below do not add up to $100\,\%$.
Refer to captionRefer to caption
Figure 2: A schematic diagram of possible key processes that drive the evolutionary phases of a triple evolution leading to the formation of a double white dwarf in the LISA frequency bandwidth. The diagram showcases key stages, including mass transfer, common envelope phases, ZLK oscillations that enhance eccentricity, and eventual binary evolution. The tertiary star plays a critical role in shaping the inner binary’s dynamics, either by inducing orbital changes or facilitating interactions that lead to the formation of the LISA double white dwarf. Blue and green circles represent the inner binary main-sequence stars. A red circle with an arrow pointing to the inner binary represents the tertiary star. A dashed arrow indicates a distant tertiary star that is too far to significantly influence the inner binary. Circles inside a brown solid ellipse show the binary. A dashed circle represents the common envelope process. A white circle represents a white dwarf, and a multi-colored circle represents a post-merger star.
Figure 3: A Venn diagram illustrating the overlap of the different evolutionary processes. Among the five processes, the Inner Binary channel is the only one that does not require mandatory assistance from the tertiary star to produce a LISA double white dwarf. In contrast, the other four processes rely on the tertiary star to bring the system into the LISA frequency bandwidth. These processes are not mutually exclusive and exhibit significant overlap.
@@ 3.1 Induced mass transfer
Figure 4: Comparison of the evolution of a triple system with and without a tertiary star. Panels (a) and (b) illustrate the evolution of the inner binary with and without the third star respectively. In the system with the tertiary star, mass transfer is induced by perturbations from the third star, allowing the system to eventually enter the LISA frequency bandwidth. When evolved without a tertiary star the binary components remain too far apart to interact. Such a system does not enter the LISA frequency bandwidth. The legends are similar to those in Figure 2. In addition, the yellow circle represents a star in the AGB phase.
@@
Figure 5: Comparative evolution of the properties of the inner binary of the triple with and without a third star. The three panels from top to bottom display the zoomed-in evolution of key parameters: eccentricity, semi-major axis, and radius of the two stars, respectively. In the case of the triple system, the inner binary experiences unstable mass transfer, causing it to shrink further and eventually enter the LISA band. Meanwhile, the inner binary when evolved without a tertiary star undergoes mass loss due to winds, resulting in an increase in its semi-major axis and a widening of the orbit. Here $a_{1}\,(\mathrm{triple})$ and $e_{1}\,(\mathrm{triple})$ represent the semi-major axis and eccentricity of the inner binary evolved with a tertiary star, while $a_{1}\,(\mathrm{binary})$ and $e_{1}\,(\mathrm{binary})$ show the semi-major axis and eccentricity of the same inner binary evolved without the tertiary star. Additionally, $r_{1}$ and $r_{2}$ represent the radii of the primary and secondary stars in the inner binary, respectively.
@@
When comparing the triple runs to the inner binary runs (where the inner binary from the same triple population evolves without the tertiary star; see Section 2). We find that approximately $9.0\,\%\pm 0.1\,\%$ of the Galactic LISA double white dwarfs undergo a mass transfer episode solely induced by the presence of a third star. These inner binaries would not have interacted over a Hubble time without the presence of the tertiary star. About $52\,\%\pm 10\,\%$ of Galactic LISA double white dwarfs initiate mass transfer at a different time due to the influence of a tertiary star. Thus, a total of 61% of systems experience induced mass transfer (see Figure 3). In contrast, the corresponding isolated binaries would either not undergo mass transfer or interact at a different time. This timing difference for the onset of mass transfer is crucial for driving the merger in the inner binary or forming a short-period inner binary that can enter the LISA frequency bandwidth within Hubble time (see the below example for more details). Also, we note that only a negligible fraction ($\sim 10^{-6}$) of systems undergo stable mass transfer without eventually becoming unstable.
@@
As examples we show the evolution of a triple system and the inner binary without the tertiary star in Figure 4. The triple system starts with $m_{1}=1.61\,\mathrm{\mathrm{M_{\odot}}}$, $m_{2}=1.28\,\mathrm{\mathrm{M_{\odot}}}$, and $m_{3}=1.16\,\mathrm{\mathrm{M_{\odot}}}$ as the tertiary star. The inner binary is initially eccentric, with an eccentricity $e_{1}=0.18$ and a semi-major axis $a_{1}=42.6\,\mathrm{au}$, while the tertiary orbit is much wider, with a semi-major axis of $a_{2}\approx 2980\,\mathrm{au}$ and an eccentricity of $e_{2}=0.26$. The initial mutual inclination between the inner and outer orbits is 99. Without the tertiary star the system evolves as an isolated binary. The relatively large initial semi-major axis allows the binary components to evolve into white dwarfs without a mass transfer episode. During the course of evolution of the binary mass loss due to winds further widens the system, as shown in Figure 5. The system continues to lose orbital energy via gravitational-wave radiation but remains too wide ($a\approx 32.1\,\rm{au}$) to enter the LISA frequency bandwidth within a Hubble time. However, with the third star present, the system undergoes significant orbital shrinkage of the inner binary during its evolution and ultimately enters the LISA frequency band. At around $2.2\,\rm{Gyr}$, the primary star of the inner binary becomes a red giant with a convective envelope. The ZLK effect, combined with tidal friction (Fabrycky & Tremaine, 2007), shrinks the inner binary’s orbit from $a_{1}\approx 42.6\,\mathrm{au}$ to $1\,\mathrm{au}$. This shortening allows the primary star to fill its Roche lobe and initiate mass transfer to its companion. The mass transfer becomes unstable due to the high mass ratio and the system enters a CE phase, which further reduces the inner binary’s orbit to $0.1\,\mathrm{au}$. This results in a $0.41\,\mathrm{\mathrm{M_{\odot}}}$ white dwarf and a $1.28\,\mathrm{\mathrm{M_{\odot}}}$ main-sequence star. Later, at around $5\,\rm{Gyr}$, the main-sequence companion evolves to an AGB star, eventually leading to a second CE phase. This produces a short-period double white dwarf with component masses of $0.41\,\mathrm{\mathrm{M_{\odot}}}$ and $0.23\,\mathrm{\mathrm{M_{\odot}}}$, which enters the LISA frequency bandwidth after a few million years. Thus, in this example, the binary would not be to become a LISA source without the tertiary.
@@ 3.2 Outer binary channel
About $16.0\,\%\pm 0.3\,\%$ of Galactic LISA double white dwarfs originate from triples where there was a merger of the inner binary stars. We identify four scenarios which lead to mergers. First, if the inner and outer orbital planes are highly inclined with respect to each other, the perturbations from the tertiary companion shrinks the orbit of the inner binary through the combination of the ZKL effect and tides. The tightened triple system has an earlier CE episode than in an isolated binary. The CE evolution results in a merger, forming a rejuvenated star bound to the former tertiary star as a binary system. This binary can evolve into the LISA frequency bandwidth. Second, in comparison to the first, the triples start out with a relatively short-period inner binary and a distant tertiary. Here, the binary comes into contact and merges without the aid of the tertiary star. The merged star further evolves with the tertiary star to enter the LISA frequency bandwidth. Third, a TCE can cause a merger of the inner binary stars (see Section 3.4 for more details). Fourth, the orbit of the inner binary can widen due to mass transfer and winds, making the entire system dynamically unstable which eventually leads a chaotic evolution of the orbits and the merger of the inner binary stars.
@@
Panel (a) of Figure 6 shows an example where the inner binary merges to form a new star. The system starts with $m_{1}=3\,\mathrm{\mathrm{M_{\odot}}}$ and $m_{2}=2.94\,\mathrm{\mathrm{M_{\odot}}}$, and $m_{3}=1.89\,\mathrm{\mathrm{M_{\odot}}}$ as the tertiary star. The inner binary is initially circular with a semi-major axis $\mathrm{a_{1}}=0.3\,\mathrm{au}$. The tertiary orbit is relatively tight, with an outer semi-major axis $\mathrm{a_{2}}=4.7\,\mathrm{au}$ and is eccentric with $\mathrm{e_{2}}=0.56$. The mutual inclination between the inner and outer orbits is $62^{\circ}$. The most massive of all three stars is in the inner binary ($m_{1}=3\,\mathrm{\mathrm{M_{\odot}}}$) which, at about $370\,\rm Myr$, initiates an unstable mass transfer episode onto the secondary resulting in the merger of the two stars. This merger leads to the formation of a new star with a mass $m_{r}=5.69\,\mathrm{\mathrm{M_{\odot}}}$. The remaining post-merger binary composed of the rejuvenated star and the tertiary companion subsequently undergoes and survives two more CE episodes: one when the rejuvenated star enters the AGB at about $390\,\rm Myr$ and another when the tertiary companion becomes an red giant star at about $1.4\,\rm Gyr$. The second CE leads to the formation of a circular double white dwarf with a semi-major axis of $9\times 10^{-3}\,\rm{au}$ later entering the LISA bandwidth after $\sim 1\,\rm{Gyr}$. An isolated binary with the same properties as the inner binary of such triple would not enter the LISA band.
Refer to captionRefer to caption
@@
Figure 6: Schematic diagrams of systems that enter the LISA frequency bandwidth after following the outer binary channel and those in which the triple eject the tertiary during the course of evolution. Panel (a) depicts an example of systems following the outer binary channel. In this scenario, the inner binary merges to form a rejuvenated star, which later enters the LISA frequency bandwidth along with the tertiary star. Panel (b) illustrates an example system that initially includes a bound third star, which facilitates a common envelope phase in the inner binary but is later ejected. The inner binary subsequently enters the LISA frequency bandwidth.The legends are similar to those in Figure 2. In addition, the yellow circle represents a star in the asymptotic giant branch (AGB) phase.

### v2

@@ 2.2 Initial conditions
Figure 1: Initial parameter distributions. The left panel shows the mass distributions ($m_{i}$), where $m_{1}$ and $m_{2}$ denote the masses of the inner binary components, and $m_{3}$ represents the mass of the tertiary. The middle panel displays the semi-major axis distributions($a_{i}$) for the inner ($a_{1}$) and outer ($a_{2}$) orbits, respectively. The right panel illustrates the eccentricity distributions($e_{i}$) of the inner ($e_{1}$) and outer ($e_{2}$) orbits, respectively. (See Sect. 2.2 for more details.)
@@
We draw the primary mass $m_{1}$ from a Kroupa initial mass function (2001MNRAS.322..231K) between $1$ and $8\,\mathrm{M_{\odot}}$. Furthermore, we follow functions from 2017ApJS..230...15M to sample the orbital period ($0.2\leq\log({T}_{1}/\text{days})\leq 8$) and secondary mass ( $0.08\,\mathrm{M_{\odot}}\leq m_{2}\leq m_{1}$) of the inner binary, and calculate the semi-major axis $a_{1}$ from Kepler’s law. Similarly, the orbital period of the outer binary also follows 2017ApJS..230...15M, where we assume that the inner binary is represented as a single star with a mass of $m_{1}+m_{2}$. We allowed the tertiary mass, $m_{3}$, to be more massive than the total mass of the inner binary in certain cases, and an extrapolated mass ratio distribution from 2017ApJS..230...15M is used to sample $m_{3}$. In addition, we sample the eccentricities of the inner ($e_{1}$) and outer ($e_{2}$) orbits from 2017ApJS..230...15M and randomly sample the spatial orientations of the inner and outer orbital frames from isotropic distributions.
We rejected any star whose radius exceeds the Roche-lobe radius on the zero-age main-sequence (1983ApJ...268..368E) and reject any system which would be dynamically unstable (2022MNRAS.516.4146V) at the start of the simulation. Any Roche lobe overflow or dynamical instabilities are modeled using prescriptions in MSE during the evolution (see Sect. 2.1 for more details).
@@ 2.3 Construction of Galactic double white dwarf population
[…]
From a population of $10^{5}$ simulated triple systems generated with MSE, we select only those that evolve into double white dwarf binaries emitting gravitational waves within the LISA sensitivity band ($10^{-4},\mathrm{Hz}$$0.1,\mathrm{Hz}$) at some point in their evolution. We use this subset to reconstruct the Galactic LISA DWD population by assigning systems to star particles in a Milky Way–like galaxy from the TNG50 simulation. Each star particle represents a population of stars spanning a range of masses, with an average mass of $\sim 10^{5}\,\mathrm{M_{\odot}}$. By populating each particle with our simulated systems, we imprint our assumed initial conditions (Sect. 2.2) and multiplicity fractions onto the TNG50 galaxy. The number of DWDs assigned to each star particle is obtained by scaling the total number of systems in the MSE sample, $N_{\textrm{DWD, MSE}}$, by the particle’s stellar mass, $M_{\star}$, at redshift $z=0$:
$N_{\textrm{DWD},\star}=\frac{N_{\textrm{DWD, MSE}}}{M_{\textrm{tot, MSE}}}\times M_{\star}.$ (1)
For each star particle, we then randomly sample $N_{\textrm{DWD},\star}$ binaries from our subset. To ensure consistency in formation history, we only retain systems whose formation times precede the particle’s age. For such systems, we account for gravitational wave radiation reaction between their formation and the particle’s age. Evolving every binary fully within MSE would be computationally prohibitive, so we adopt a hybrid approach. For circular binaries, the tertiary companion has no appreciable effect once the system enters the LISA band. These binaries are therefore evolved analytically forward in time, using the gravitational wave emission formula of 1964PhRv..136.1224P, until they reach the age of the host particle. In contrast, eccentric binaries (much rarer but significantly influenced by the tertiary) are included only if their simulated formation times lie within $\pm 100{,}000$ years of the particle’s age (a choice set mainly by computational constraints; see Sect. 4.1.2).
The total stellar mass of the simulated population is given by
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{for }f_{t}\neq 0,$
@@ 3 Key processes that shape the triple evolutionary pathways
In this section we present a brief overview of the five key processes characteristic of triple evolution that shape the evolutionary pathways leading to the formation of LISA double white dwarfs:
@@
    Triple common envelope: When the tertiary star overflows its Roche lobe, mass transfer onto the inner binary initiates a triple common envelope phase, which further drives the formation of a LISA double white dwarf.
@@
We show a schematic diagram for primary processes that drive evolutionary pathways leading to LISA double white dwarfs from triples in Fig. 2. In the following subsections, we discuss these processes in greater depth with detailed examples. All uncertainties presented below are estimated by scaling up the fractional Poisson error from the intrinsic population evolved with MSE. Figure 3 shows that the processes are not mutually exclusive. For instance, about 4.8% of Galactic LISA double white dwarfs from triples undergo a TCE phase and induced mass transfer, leading to a merger in the inner binary. Hence, the relative percentages quoted below do not add up to $100\,\%$.
Figure 2: Diagram of possible key processes that drive the evolutionary phases of a triple evolution leading to the formation of a double white dwarf in the LISA frequency bandwidth. The diagram showcases key stages, including mass transfer, common envelope phases, ZLK oscillations that enhance eccentricity, and eventual binary evolution. The tertiary star plays a critical role in shaping the inner binary’s dynamics, either by inducing orbital changes or facilitating interactions that lead to the formation of the LISA double white dwarf. The circles represent the index of the star, with blue, green, and red indicating the primary, secondary, and tertiary stars, respectively. The filling inside each circle represents the star’s evolutionary phase: purple for the main-sequence and white for a white dwarf. A dashed arrow denotes a distant tertiary star that is too far to significantly influence the inner binary. A multi-colored circle represents a post-merger star, with the two colors signifying the components that have merged.
Figure 3: Venn diagram illustrating the overlap of the different evolutionary processes. Among the five processes, the inner binary channel is the only one that does not require assistance from the tertiary star to produce a LISA double white dwarf. In contrast, the other four processes rely on the tertiary star to bring the system into the LISA frequency bandwidth. These processes are not mutually exclusive and exhibit significant overlap.
@@ 3.1 Induced mass transfer
Figure 4: Comparison of the evolution of a triple system with and without a tertiary star. Panels (a) and (b) illustrate the evolution of the inner binary with and without the third star respectively. In the system with the tertiary star, mass transfer is induced by perturbations from the third star, allowing the system to eventually enter the LISA frequency bandwidth. When evolved without a tertiary star the binary components remain too far apart to interact. Such a system does not enter the LISA frequency bandwidth. The legends are similar to those in Fig. 2. In addition, the yellow filling represents a star in the Asymptotic giant branch phase.
@@
Figure 5: Comparative evolution of the properties of the inner binary of the triple with and without a third star. The three panels, from top to bottom, display the zoomed-in evolution of key parameters of the two stars: eccentricity, semi-major axis, and radius. In the case of the triple system, the inner binary experiences unstable mass transfer, causing it to shrink further and eventually enter the LISA band. Meanwhile, the inner binary when evolved without a tertiary star undergoes mass loss due to winds, resulting in an increase in its semi-major axis and a widening of the orbit. Here $a_{1}\,(\mathrm{triple})$ and $e_{1}\,(\mathrm{triple})$ represent the semi-major axis and eccentricity of the inner binary evolved with a tertiary star, while ${a}\,({\rm binary})$ and ${e}\,({\rm binary})$ show the semi-major axis and eccentricity of the same inner binary evolved without the tertiary star. Additionally, $r_{1}$ and $r_{2}$ represent the radii of the primary and secondary stars in the inner binary, respectively.
@@
When comparing the triple runs to the inner binary runs (where the inner binary from the same triple population evolves without the tertiary star; see Sect. 2). We find that approximately $9^{+2}_{-3}\,\%$ of the Galactic LISA double white dwarfs undergo a mass transfer episode solely induced by the influence of a third star. These inner binaries would not have interacted over a Hubble time without the presence of the tertiary star. About $52^{+6}_{-7}\,\%$ of Galactic LISA double white dwarfs initiate mass transfer at a different time due to the influence of a tertiary star. Thus, a total of $61^{+8}_{-10}\,\%$ of systems experience induced mass transfer (see Fig. 3). The mass transfer is induced due to the combined effect of dynamical stability constraints and dynamical interactions resulting from the third star. It is not trivial to quantify the precise influence of the third star. This mass transfer is often found to have occurred at various evolutionary stages well before the formation of double white dwarfs. In contrast, the corresponding isolated binaries would either not undergo mass transfer or interact at a different time. This timing difference for the onset of mass transfer is crucial for driving the merger in the inner binary or forming a short-period inner binary that can enter the LISA frequency bandwidth within Hubble time (see the example below for more details). Also, we note that only a negligible fraction ($\sim 10^{-6}$) of systems undergo stable mass transfer without eventually becoming unstable.
@@
As examples we show the evolution of a triple system and the inner binary without the tertiary star in Fig. 4. The triple system starts with $m_{1}=1.61\,\mathrm{\mathrm{M_{\odot}}}$, $m_{2}=1.28\,\mathrm{\mathrm{M_{\odot}}}$, and $m_{3}=1.16\,\mathrm{\mathrm{M_{\odot}}}$ as the tertiary star. The inner binary is initially eccentric, with an eccentricity $e_{1}=0.18$ and a semi-major axis $a_{1}=42.6\,\mathrm{au}$, while the tertiary orbit is much wider, with a semi-major axis of $a_{2}\approx 2980\,\mathrm{au}$ and an eccentricity of $e_{2}=0.26$. The initial mutual inclination between the inner and outer orbits is $9$. Without the tertiary star the system evolves as an isolated binary. The large initial semi-major axis allows the binary components to evolve into white dwarfs without a mass transfer episode. During the course of evolution of the binary mass loss due to winds further widens the system, as shown in Fig. 5. The system continues to lose orbital energy via gravitational wave radiation but remains too wide ($a\approx 32.1\,\rm{au}$) to enter the LISA frequency bandwidth within a Hubble time. However, with the third star present, the system undergoes significant orbital shrinkage of the inner binary during its evolution and ultimately enters the LISA frequency band. At around $2.2\,\rm{Gyr}$, the primary star of the inner binary becomes a red giant with a convective envelope. The ZLK effect, combined with tidal friction (2007ApJ...669.1298F), shrinks the inner binary’s orbit from $a_{1}\approx 42.6\,\mathrm{au}$ to $1\,\mathrm{au}$. This shortening allows the primary star to fill its Roche lobe and initiate mass transfer to its companion. The mass transfer becomes unstable due to the high mass ratio and the system enters a CE phase, which further reduces the inner binary’s orbit to $0.1\,\mathrm{au}$. This results in a $0.41\,\mathrm{\mathrm{M_{\odot}}}$ white dwarf and a $1.28\,\mathrm{\mathrm{M_{\odot}}}$ main-sequence star. Later, at around $5\,\rm{Gyr}$, the main-sequence companion evolves to an AGB star, eventually leading to a second CE phase. This produces a short-period double white dwarf with component masses of $0.41\,\mathrm{\mathrm{M_{\odot}}}$ and $0.23\,\mathrm{\mathrm{M_{\odot}}}$, which enters the LISA frequency bandwidth after a few million years. Thus, in this example, the binary would not be to become a LISA source without the tertiary.
@@ 3.2 Outer binary channel
About $16^{+3}_{-3}\,\%$ of Galactic LISA double white dwarfs originate from triples where there was a merger of the inner binary stars. We identified four scenarios which lead to mergers. First, if the inner and outer orbital planes are highly inclined with respect to each other, the perturbations from the tertiary companion shrinks the orbit of the inner binary through the combination of the ZKL effect and tides. The tightened triple system has an earlier CE episode than in an isolated binary. The CE evolution results in a merger, forming a rejuvenated star bound to the former tertiary star as a binary system. This binary can evolve into the LISA frequency bandwidth. Second, in comparison to the first, the triples start out with a short-period inner binary and a distant tertiary. Here, the binary comes into contact and merges without the aid of the tertiary star. The merged star further evolves with the tertiary star to enter the LISA frequency bandwidth. Third, a TCE can cause a merger of the inner binary stars (see Sect. 3.4 for more details). Fourth, the orbit of the inner binary can widen due to mass transfer and winds, making the entire system dynamically unstable which eventually leads a chaotic evolution of the orbits and the merger of the inner binary stars.
@@
Panel (a) of Fig. 6 shows an example where the inner binary merges to form a new star. The system starts with $m_{1}=3\,\mathrm{\mathrm{M_{\odot}}}$ and $m_{2}=2.94\,\mathrm{\mathrm{M_{\odot}}}$, and $m_{3}=1.89\,\mathrm{\mathrm{M_{\odot}}}$ as the tertiary star. The inner binary is initially circular with a semi-major axis $\mathrm{a_{1}}=0.3\,\mathrm{au}$. The tertiary orbit is tight, with an outer semi-major axis $\mathrm{a_{2}}=4.7\,\mathrm{au}$ and is eccentric with $\mathrm{e_{2}}=0.56$. The mutual inclination between the inner and outer orbits is $62^{\circ}$. The most massive of all three stars is in the inner binary ($m_{1}=3\,\mathrm{\mathrm{M_{\odot}}}$) which, at about $370\,\rm Myr$, initiates an unstable mass transfer episode onto the secondary resulting in the merger of the two stars. This merger leads to the formation of a new star with a mass $m_{r}=5.69\,\mathrm{\mathrm{M_{\odot}}}$. The remaining post-merger binary composed of the rejuvenated star and the tertiary companion subsequently undergoes and survives two more CE episodes: one when the rejuvenated star enters the AGB at about $390\,\rm Myr$ and another when the tertiary companion becomes an red giant star at about $1.4\,\rm Gyr$. The second CE leads to the formation of a circular double white dwarf with a semi-major axis of $9\times 10^{-3}\,\rm{au}$ later entering the LISA bandwidth after $\sim 1\,\rm{Gyr}$. An isolated binary with the same properties as the inner binary of such a triple would not enter the LISA band.
@@
Figure 6: Schematic diagrams of systems that enter the LISA frequency bandwidth after following the outer binary channel and those in which the triple eject the tertiary during the course of evolution. Panel (a) depicts an example of systems following the outer binary channel. In this scenario, the inner binary merges to form a rejuvenated star, which later enters the LISA frequency bandwidth along with the tertiary star. Panel (b) illustrates an example system that initially includes a bound third star, which facilitates a common envelope phase in the inner binary but is later ejected. The inner binary subsequently enters the LISA frequency bandwidth.The legends are similar to those in Fig. 2. In addition, the yellow circle represents a star in the asymptotic giant branch phase.